Good day.

How to solve below question?
5. 7/8 ÷ 21/4 = ___________?

Answer by Frosty7(4)

7/8 ÷ 21/4
= 7/8 x 4/21 (4/21 is the reciprocal of 21/4)
= 28/168
= 1/6
Hope this helps!